If sales price remained the same and variable costs increased the contribution margin would
Increase
Decrease
Remain the same
Answer #1

If sale price same and variable cost increase then contribution margin would be decrease.

Example if sales 10 and variable cost 5 then contribution margin = (10-5) = 5 per unit

ANd if sales 10 and variable cost 6 then contribution margin (10-6) = 4 per unit

so answer is b) Decrease